OpenStudy (anonymous):

How does one do this problem? Simplify. Assume that no denominator is equal to zero. (4g^2h^3)^3

OpenStudy (anonymous):

Please.. help.

OpenStudy (anonymous):

Lets start off with something easy. \[(x^6)^2\] Do you know how to use rules of exponents to "simplify" this?

OpenStudy (anonymous):

well first you need to know how to multiply powersx power

OpenStudy (anonymous):

@agentc0re Would it be x^12

OpenStudy (anonymous):

Not quite. Lets think of something a bit more basic. What is \[6^2\]?

OpenStudy (anonymous):

36. :P

OpenStudy (anonymous):

.....

OpenStudy (anonymous):

Sorry my browser just crapped out and i'm gonna have to retype my answer.

OpenStudy (anonymous):

what we can see is that we can group things first to obtain the actual answer. so in my simple example of\[(x^6)^2 = x^{(6^2)} = x^{36}\] Now, all we have to do is do this with your problem. Let me try and help you group everything together. \[{(4g^{(2h^3)})}^3 = {4^3g^{3(2h)^{3*3}}}\] Remember your rules of arithmetic. You have to do powers before multiplication. Let me know what you get and we'll see if it's right or not and i can help better explain if need be. :D